Written in standard form an equation of the line passing through the given points with the given slope. slope=1/6;(1,-5)
Answer by lwsshak3(11628)

Written in standard form an equation of the line passing through the given points with the given slope. slope=1/6;(1,-5)
***
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
For given problem:
equation: y=x/6+b
solve for b using coordinates of given point (1,-5)
-5=1/6+b
b=-5-1/6=-31/6
equation: y=x/6-31/6